# Hyman et al. 2025

### v2

@@ 1 Introduction
Both pre- and in-season harvest projections are typically estimated using historical, temporally indexed harvest information from previous years (carruthers2014evaluating, ). Recreational fishing season closure dates are often established before the season begins to provide private anglers and for-hire businesses with a clear expectation of the total season duration. This approach relies on projections for when the recreational quota will be met before any harvest data from the present year are available. (farmer2015forecasting, ; farmer2020forecasting, ). Additionally, temporally indexed recreational harvest estimates are seldom available on timelines required for in-season management (carter2015nowcasting, ; MRIPData2021, ). Due to the substantial number of participants in many recreational fisheries, recreational harvest is often estimated using sampling data from dockside interviews and mail surveys of effort, followed by survey-based extrapolation methods (MRIPData2021, ; MRIP2023, ). The time required to implement this methodology, especially the mail survey, results in estimate reporting lags of up to several months after sampling. Consequently, managers are often forced to rely on limited in-season and past harvest information to formulate and update harvest forecasts as the season progresses.
@@
Unfortunately, past harvest data can be an unreliable predictor of future harvest rates when regulations change (carter2015nowcasting, ). Angler behaviors, including decisions about whether to fish and which species to target, are heavily influenced by fishing regulations (farmer2020forecasting, ; Hyman2024Effort, ). Anglers often focus on popular species that are currently open for harvest, and may concentrate their efforts on a particular species if its season is shortened (abbott2018status, ; trudeau2022lower, ). When harvests exceed projections based on past estimates, managers may need to implement post-hoc accountability measures, such as reducing the quota for the subsequent year (i.e., a payback; methot2014implementing, ). Participants in the recreational fishery may perceive these measures as punitive, despite their adherence to regulations, potentially leading to decreased future compliance and undermining management integrity (cowan2011faith, ; cowan2011red, ). Therefore, techniques that bridge the gap between harvest data availability and real-time management needs are valuable tools for ensuring sustainable fisheries and promoting recreational angler satisfaction.
@@
Modeling frameworks leveraging relationships between readily available social, environmental, and management variables and harvest have demonstrable success in improving the accuracy of both pre-season projection and in-season monitoring estimates. For example, a suite of sophisticated modeling approaches has been applied to reliably forecast landings from a broad range of species, including Gulf of Mexico (Gulf) vermilion snapper (Rhomboplites aurorubens) and gray snapper (Lutjanus griseus) (farmer2015forecasting, ), Gulf red snapper (Lutjanus campechanus)(farmer2020forecasting, ), Atlantic menhaden (Brevoortia tyrannus)(hanson2006forecasting, ), Pacific giant octopus (Enteroctopus dofleini) (nagano2023predicting, ), and the Indian oil sardine (Sardinella longiceps) (holmes2021improving, ). Such methods can be used to set season durations (farmer2020forecasting, ) or ‘nowcast’ in-season catch rates (carter2015nowcasting, ) using combinations of fishing regulations, socioeconomic factors, weather conditions, and internet search traffic. In fact, these modeling approaches can be considerable improvements to more typical projections primarily based on past harvest data (carter2015nowcasting, ).
@@ 2.1 Data sources
We compiled recreational gag harvest data for the for-hire sector from the National Marine Fisheries Service (NMFS) Marine Recreational Information Program (MRIP) For-Hire Telephone Survey (FHTS) combined with data from the Angler Point Access Intercept Survey (APAIS) and private recreational data from the Florida Fish and Wildlife (FWC) State Reef Fish Survey (SRFS). This approach aligns with the data sources currently used for gag management, as both the GMFMC and NOAA NMFS monitor recreational gag landings using SRFS and FHTS/APAIS data.
@@ 2.3 Analysis
We used the Stan probabilistic programming language for Bayesian modeling (gelman2015stan, ), accessed via R (Rcore2024, ), to fit all models. Each model involved four parallel Markov chains with 5,000 iterations for warm-up and another 5,000 for posterior sampling, yielding 20,000 total draws for inference. Covariates were deemed significant if their 80% confidence intervals excluded 0 (ipccAR6, ; kruschke2021bayesian, ; van2021bayesian, ). All graphics were produced using the ggplot2 package in R (ggplot2, ).
@@ 3.1 Model selection and diagnostics
Time-series plots comparing observed and predicted values, along with posterior predictive checks, indicated that model $g_{2}$ provided a good fit to the observed data. The posterior predictive median and 80% CI of monthly estimates aligned closely with in-sample estimates of monthly gag harvest rates (Fig. 1a). Additionally, diagnostic scatter-plots of median predicted versus observed gag harvest rates and comparisons of empirical distributions to posterior predictive distributions derived from model $g_{2}$ demonstrated satisfactory model performance (Figs. 1b and 1c).

### v3

@@ 1 Introduction
Both pre- and in-season harvest projections are typically estimated using historical, temporally indexed harvest information from previous years (carruthers2014evaluating, ). Recreational fishing season closure dates are often established before the season begins to provide private anglers and for-hire businesses with a clear expectation of the total season duration. This approach relies on projections for when the recreational quota will be met before any harvest data from the present year are available (farmer2015forecasting, ; farmer2020forecasting, ). Additionally, temporally indexed recreational harvest estimates are seldom available on timelines required for in-season management (carter2015nowcasting, ; MRIPData2021, ). Due to the substantial number of participants in many recreational fisheries, recreational harvest is often estimated using sampling data from dockside interviews and mail surveys of effort, followed by survey-based extrapolation methods (MRIPData2021, ; MRIP2023, ). The time required to implement this methodology, especially the mail survey, results in estimate reporting lags of up to several months after sampling. Consequently, managers are often forced to rely on limited in-season and past harvest information to formulate and update harvest forecasts as the season progresses.
@@
Unfortunately, past harvest data can be an unreliable predictor of future harvest rates when regulations change (carter2015nowcasting, ). Angler behaviors, including decisions about whether to fish and which species to target, are heavily influenced by fishing regulations (farmer2020forecasting, ; Hyman2024Effort, ). Anglers often focus on popular species that are currently open for harvest and may concentrate their efforts on a particular species if its season is shortened (abbott2018status, ; trudeau2022lower, ). When harvests exceed projections based on past estimates, managers may need to implement post-hoc accountability measures, such as reducing the quota for the subsequent year (i.e., a payback; methot2014implementing, ). Participants in the recreational fishery may perceive these measures as punitive, despite their adherence to regulations, potentially leading to decreased future compliance and undermining management integrity (cowan2011faith, ; cowan2011red, ). Therefore, techniques that bridge the gap between harvest data availability and real-time management needs are valuable tools for ensuring sustainable fisheries and promoting recreational angler satisfaction.
@@
Modeling frameworks leveraging relationships between readily available social, environmental, and management variables and harvest have demonstrable success in improving the accuracy of both pre-season projection and in-season monitoring estimates. For example, a suite of sophisticated modeling approaches has been applied to reliably forecast landings from a broad range of species, including Gulf of Mexico (Gulf) vermilion snapper (Rhomboplites aurorubens) and gray snapper (Lutjanus griseus) (farmer2015forecasting, ), Gulf red snapper (Lutjanus campechanus)(farmer2020forecasting, ), Atlantic menhaden (Brevoortia tyrannus)(hanson2006forecasting, ), Pacific giant octopus (Enteroctopus dofleini) (nagano2023predicting, ), and the Indian oil sardine (Sardinella longiceps) (holmes2021improving, ). Such methods can be used to set season durations (farmer2020forecasting, ) or ‘nowcast’ in-season catch rates (carter2015nowcasting, ) using combinations of fishing regulations, socioeconomic factors, weather conditions, and internet search traffic. In fact, these modeling approaches can represent considerable improvements to more typical projections primarily based on past harvest data (carter2015nowcasting, ).
@@ 2.1 Data sources
We compiled recreational gag harvest data for the for-hire sector from the National Marine Fisheries Service (NMFS) Marine Recreational Information Program (MRIP) For-Hire Survey (FHS) combined with data from the Angler Point Access Intercept Survey (APAIS) and private recreational data from the Florida Fish and Wildlife (FWC) State Reef Fish Survey (SRFS). This approach aligns with the data sources currently used for gag management, as both the GMFMC and NOAA NMFS monitor recreational gag landings using SRFS and FHS/APAIS data.
@@ 2.3 Analysis
We used the Stan probabilistic programming language for Bayesian modeling (gelman2015stan, ), accessed via R (Rcore2024, ), to fit all models. Each model involved four parallel Markov chains with 5,000 iterations for warm-up and another 5,000 for posterior sampling, yielding 20,000 total draws for inference. Covariates were deemed significant if their 80% confidence intervals excluded 0 (ipccAR6, ; kruschke2021bayesian, ; van2021bayesian, ). All graphics and auxiliary analyses were produced using the ggplot2 package in R (ggplot2, ).
@@ 3.1 Model selection and diagnostics
Time-series plots comparing observed and predicted values, along with posterior predictive checks, indicated that model $g_{2}$ provided a good fit to the observed data. The posterior predictive median and 80% CI of monthly estimates aligned closely with in-sample estimates of monthly gag harvest rates (Fig. 1a). Additionally, diagnostics derived from model $g_{2}$ demonstrated satisfactory model performance (Figs. 1b and 1c).